True Basic

A sketch of John Kemeny for the Dartmouth Alumni Magazine

Nardi Reeder Campion

December 13, 2001

The newspaper said John G. Kemeny, 13th president of Dartmouth College, died of heart failure. Clearly this was a mistake. John Kemeny’s heart never failed anyone.

It was altogether fitting that at St. Thomas Episcopal Church we prayed for ”the eternal rest of President Kemeny’s immortal soul.” Although he was a Jew, in John Kemeny there was no east nor west, no north nor south. He was an ecumenical man in the truest sense of the word: ”of or from the whole world.”

During that service Gil Tanis ’38, who served President Kemeny as an executive officer, rose to speak. The laudatory obits in the Times and the Globe, Tanis said, ”ignored an important fact about Kemeny: his love of teaching. No matter how crowded his schedule, John Kemeny found time to teach. Teaching was his joy.”

In fact, John Kemeny agreed to become president of Dartmouth in 1970 on one condition: that he be allowed to continue teaching. When the chairman of the Board of Trustees objectedthe presidency was too demanding-Kemeny asked what the Trustees would say if he wanted two hours off each week to play golf. That did it.

Throughout his presidency he taught two classes a year, from far-out computer science to freshman math. Having lamented that ”math is the only subject you can study for l4 years and not learn a single thing that has been done since 1800,” he and his math colleague, Laurie Snell, strung logic, probability, and matrix algebra into a course so fresh that it merited a new name: Finite Math. Professor Kemeny was ”remarkable in his ability to understand what a student didn’t understand and re-explain it,” recalls former student Rob Salzman ’76.

I once experienced John’s brilliant teaching. In 1978, my husband and I attended a Dartmouth Alumni College session on ”Where Have All The Heroes Gone?” The director, Professor Jim Epperson, had invited President Kemeny to speak on his hero, Albert Einstein. John began by saying that Einstein, a shy, humble man who never performed an experiment, had, by just sitting in a room and thinking, changed the entire universe. Then Kemeny went to the blackboard and with breathtaking clarity taught 300 older people the general theory of relativity, using an example of two elevators. He finished with a flourish, writing on the board E=MC2. We jumped to our feet and gave him an ovation. Effortlessly, John Kemeny had carried us beyond our depth into the elegant world of pure mathematics.

Perhaps because he was excruciatingly shy himself, John was ever mindful of the needs of others. Before Jim Freedman became president of Dartmouth, Kemeny asked to see Sheba Freedman alone. He talked with her about the difficulties she would face as Dartmouth’s first first lady to be on the faculty. ”John realized,” Sheba says, ”that if I was ambitious enough to get my Ph.D. in psychology, teaching was important to me. How smart he was-and how understanding.”

That was evident from the start. Lu Martin, special assistant to President Kemeny, recalls how he handled the first crisis of his administration. Two months after he took office, the bombings in Cambodia and the killings at Kent State ignited rage and fear on campuses across the nation. Kemeny rose to the challenge. By canceling classes and leading the college in a week of soul-searching and mourning, he averted a student strike.

The Manchester Union Leader denounced him in a front-page editorial headlined: ”Dartmouth Has Bought Another Lemon.” Lemon T-shirts instantly blossomed on campus; students delivered bushels of lemons to the Union Leader; and faculty presented John with a live lemon tree. At the end of the week, John was to address a rally in the field house. His wife, Jean, advised him, ”If you’re handed a lemon, make lemonade,” and she gave him a lot of lemons. President Kemeny concluded his talk by tossing lemons to an enthusiastic crowd. Afterward, lemons autographed by the president became collectors’ items.

Like everyone, John Kemeny had a few shortcomings. He did not suffer fools gladly. He did not moderate his strong opinions or his explosive temper. He did not hide his light under a bushel. He was good and he knew it. To a colleague who suggested that another administrator may have done some things better than John did, Kemeny snapped, ”Name one.”

But Kemeny could listen. ”I had the sense that when I was with him he was paying attention and cared very much about what the students had to say,” says ltob Salzman of his days as a student intern in the president’s office. ”Students were surprised by that. When they had that kind of personal contact they really had the sense he was viewing what you said as very important.”

Kemeny was not programmed for a college president’s hours. He never arrived on campus before late morning. A night person, he functioned best between midday and midnight. If concentrating on a problem, he never stopped just because the hours grew small. Once Provost Leonard Rieser ’44 ran into him at 9 a.m. ”John!” he exclaimed. ”What are you doing on campus at this hour?” ”Going home,” John replied.

Nor was John a Big Green outdoorsman. ”My idea of roughing it,” he said, ”is a hotel room with only a shower.” A fellow mathematician told meJohn once played second base on the math department softball team: ”He was as inept at base running as he was ept at everything else.”

Still, he loved Dartmouth sports. He and his math colleagues always sat together at football games, Kemeny in the middle. They made all manner of bets based on probabilities of obscure game occurrences, such as the odds on three flags being thrown twice in a quarter.

Kemeny knew so much about so many things. He began life in Hungary, a prodigy whose boyhood was uprooted by the Nazis. In 1940 his father saw what was coming for European Jews and took the family to America; Kemeny’s grandfather refused to leave and perished in the Holocaust, as did an aunt and uncle. Although John had never heard a word of English before he was 13, he graduated from Princeton in three years with enough credits to major in philosophy as well as math. When he took a year off to work on the Manhattan Project in Los Alamos, his boss was the Nobel Prize-winnerer Richard Feynman. At age 22, a year before he earned his Ph.D., Kemeny became Albert Einstein’s mathematical assistant. I once asked him why Dr. Einstein, of all people, needed a mathematician. With that gentle, mustached smile, John said, ”Einstein wasn’t very good at math.”

According to his colleagues, Kemeny’s keen analytical ability to assess needs and uncover potentialities was a prime element in his value to the College. He built the Dartmouth math department into a national model. He wrote or co-authored 13 books on math and philosophy. He and fellow Dartmouth mathematician Tom Kurtz revolutionized computing with the world’s first time-sharing system. Then they co-authored BASIC, the internationally used computer language that opened computing to a much wider public. And Kemeny presided over the most profound changes at Dartmouth in modern times: co-education, year-round scheduling, the renewal of Native American education, and active recruitment of minority students.

No problem seemed unsolvable. President Carter in 1979 selected John Kemelly to head the Three Mile Island investigation. Kemeny was horrified by the conditions he found there. But as his friend Laurie Snell puts it, ”John was a born optimist. He was always sure things could be improved.”

Even the most brilliant thinkers can be absent-minded at times. Leonard Rieser says John’s extraordinary intensity of concentration was the cause of his apparent absentmindedness. ”One time John and I were driving to Yale, where he was to speak at a luncheon. His topic was the new doctor of arts program in math, and he was giving that his full attention. Finally I said, ’John, I think we just passed Greenwich.’

”He wheeled around, and we raced north toward New Haven. John was proud of his red car with a souped-up engine, but it did require gas. He was still expounding on his topic when the tank went dry. After we got gas, the car overheated. Of course we missed the luncheon. As we walked into the room, Kingman Brewster was saying, ’Our next speaker will be John Kemeny.”’ Kemeny strolled to the podium and picked up without missing a beat.

Sometimes John’s playfulness caught people off guard. One of his vice presidents, Rod Morgan, returned from a Caribbean vacation to find this letter on his desk:

       DEAR ROD:
       It was very nice of you to send that lovely post card to tell us
       about the sunshine and your hammocle. The campus is covered
       by sleet, it is unattractive, it is horrible. The weather bureau
       is predicting the worst blizzard we have had in years for
       tomorrow. You're fired!
       Yours truly,
       JOHN

       PS. On second thought, I forgive you. We're just jealous! Welcome home.

John Kemeny’s mind retained most everything but not names. That was Jean’s department; his dependency on his wife was total. They met when she was a freshman at Smith and he, with his new Ph.D., an assistant professor of philosophy at Princeton. He was 23, she l 9. Jean did not return for her sophomore year. They were married in November 1950. Their 42-year marriage was a model of devotion. ”His loyalty is something I can absolutely depend on,” Jean wrote in her book, It’s Different at Dartmouth. And in a way, the marriage was a model for Kemeny’s relationship to Dartmouth. ”My commitment to this College is the same as my commitment to my wife: ’til death do us part,” he remarked on the tenth anniversary of his presidency. That his daughter Jennifer ’76 and son Robert ’77 chose to attend Dartmouth says something about the bonds that tied the Kemenys to one another and to the College.

After John’s sudden death, Jean wanted to talk about him. ”People didn’t know he was a real softy. Every Christmas we watched Tile Miracle On 34th Street, and every Christmas he cried. Because we loved a movie called The Bear, our daughter gave us a bear that John kept on the bedside table. Each night he said, ’Good night, Bear.’ Bear would appear regularly on my pillow wearing glasses, or a hat, or smoking a pipe.

”John was a funny and very touching person who communicated better with animals than with people. Our fluffy black cat, Who, followed him like a dog. For 15 years, John has been feeding our raccoons. [The Kemenys’ house is on a remote hill in Etna.] Now 35 raccoons come every day to the feeding station on our deck, 16 feet above the ground, to eat and nurse their babies. John would greet each, one by one.

”He liked science fiction, football games, shrimp, all kinds of puzzles, Agatha Christie, and solitude (for two). He did not enjoy socializing. Before he retired, John recognized only two flowers, the tulip and the rose, and two pieces of music, the 1812 Overture and Poor Little Buttercup. These last years he had time to enjoy Mozart, wildflowers, pileated woodpeckers, eclipses. Sometimes he liked just to sit still and think. John wasn’t a cook, but when I yearned for a Grand Marnier souffle, he collated all our cookbooks and created a perfect souffle with creme anglaise sauce, just to please me.

”The night before he died we listened to Frederika Von Staade and Kathleen Battle’s Christmas concert. He died the day after Christmas. It all happened so fast. The nausea, the awful pain-and then he just fell. I called 911 and a policeman came.

”When we got him to the hospital, he had no blood pressure. He was only 66. Jim Strickler, former dean of the Dartmouth Medical School, told me that John’s coronary artery was calcified, but in spite of all the smoking, his lungs were clear. John and I were two cultures, two countries-a Yankee from Maine and a Jew from Budapest-but,” Jean’s voice broke slightly, ”it worked.”

John G. Kemeny possessed the inner calm that comes from having decided what is important and what is not. When he presided at his last commencement, he presented a diploma to a Dartmouth Review editor and that young man handed him a piece of paper. On it he had written: ”F*** you.” President Kemeny did not turn a hair.

He signaled the dean to raise the senior class. Then, in his soft Hungarian voice, John Kemelly gave his charge to the graduates: ”The most dangerous voice you will ever hear is the evil voice of prejudice that divides black from white, man from woman, Jew from gentile. Listen to the voice that says man can live in harmony. Use your very considerable talents to make the world better.”

President John Kemeny, the ecumenical man, ended his last commencement exercises with the words he used to conclude every conmmencement, words that came from his heart: ”Women and men of Dartmouth, all mankind is your brother-and you are your brother’s keeper.”